The following measurements were recorded for the drying time, in hours, of a certain brand of latex paint:
3.4 2.5 4.8 2.9 3.6
2.8 3.3 5.6 3.7 2.8
4.4 4.0 5.2 3.0 4.8
Assuming that the measurements represent a random sample from a normal population, find a 95% prediction interval for the drying time for the next trial of
the paint
1
Expert's answer
2020-05-19T19:06:54-0400

"2.5, 2.8, 2.8, 2.9, 3.0,"

"3.3, 3.4, 3.6, 3.7, 4.0,"

"4.4, 4.8, 4.8, 5.2, 5.6"



"\sum_ix_i=2.5+ 2.8+ 2.8+2.9+3.0+""+3.3+3.4+3.6+ 3.7+ 4.0+4.4+""+4.8+4.8+5.2+5.6=56.8"

"mean=\bar{x}={\sum_ix_i\over n}={56.8\over 15}\approx3.7867"

"\sum_i(x_i-\bar{x})^2={2969.4\over 225}\approx13.1973"

"s^2={\sum_i(x_i-\bar{x})^2\over n-1}={2969.4\over 225(15-1)}={212.1\over 225}"

"s=\sqrt{s^2}=\sqrt{{212.1\over 225}}\approx0.9709"

The provided sample mean is "\bar{x}=3.7867" and the sample standard deviation is "s=0.9709."

The size of the sample is "n=15" and the required confidence level is 95%.

The number of degrees of freedom are "df=n-1=15-1=14," and the significance level is 

"\alpha=0.05."

Based on the provided information, the critical t-value for "\alpha=0.05" and "df=14" degrees of freedom is "t_c=2.144787"

The 95% confidence for the population "\mu" is computed using the following expression

"CI=(\bar{x}-t_c\times s \sqrt{1+1/n},\bar{x}+t_c\times s \sqrt{1+1/n})"

"=(3.7867-2.144787\times{0.9709\sqrt{14}\over \sqrt{15}},3.7867+2.144787\times{0.9709\sqrt{14}\over \sqrt{15}})"

"=(1.7749, 5.7985)"